OpenStudy (anonymous):

For what values of x in [0, 2*pi] does the graph of y={cos(x)}/{2+sin(x)} have a horizontal tangent? List the smaller value of x first.

OpenStudy (aum):

Horizontal tangent implies a line parallel to the x-axis which has a slope of zero. The derivative gives the slope. Find the derivative, equate it to zero and solve for x in the interval [0, 2pi].

OpenStudy (anonymous):

okay so the derivative is y'(x) = -(2sinx+1)/(sinx+2)^2

OpenStudy (freckles):

looks good i bet you used the pythagorean identity cos^2(x)+sin^2(x)=1 to clean it up into that.

OpenStudy (freckles):

so as @aum was saying you need find when y'=0

OpenStudy (freckles):

so when does 2sin(x)+1=0

OpenStudy (freckles):

You will find the unit circle most helpful here.

OpenStudy (anonymous):

so if I isolate the x it would be sin^-1(1/2) = x. What am I looking for in the unit circle?

OpenStudy (anonymous):

So is it sqrt(3)/2 and the negative of that?

OpenStudy (freckles):

2sin(x)+1=0 sin(x)=-1/2 There should be two values such that sin(x) is -1/2

OpenStudy (aum):

sin(x) = -1/2. From the unit circle, the reference angle is pi/6 radians as sin(pi/6) = 1/2 Sine is negative in the third and fourth quadrant. Therefore the angles are pi + pi/6 = 7pi/6 and 2pi - pi/6 = 11pi/6

OpenStudy (anonymous):

Thanks.